Solve the absolute value equation: | 3 – (3/4)x| = ¼
Answer by tutorcecilia(2144)

| 3 – (3/4)x| = ¼ [Change the fractions to decimals]
| 3 – .75x| = .25 [Abosolute values may have two answers]
.
3 – (.75)x = .25 [If x>0]
3-3-(.75)x = .25-3 [Solve for x]
-.75x = -2.75
-.75x/-.75 = -2.75/-.75
x = 3.666
.
3 – (.75)x = -.25 [If x<0]
3-3-.75x = -.25-3 [Solve for x]
-.75x = -3.25
-.75x/-.75 = -3.25/-.75
x=4.333
.
So x = 3.666 or x = 4.333
.
Check by plugging in each value of "x" back into the original equation.
.